OpenStudy (2yoututu11):

What is the percent of increase from 50 to 90? A. 40% B. 50% C. 80% D. 90%

OpenStudy (mathmagician):

50-100% 90-x x=90*100/50=180% So, the increase is 180%-100%=80%

OpenStudy (2yoututu11):